Last year, Evelina invested $10,000.00, part at 6% annual interest and part at 8% annual interest. If she received $760.00 in interest at the end of the year, how much did she invest at each rate?
Found 2 solutions by rfer, mananth:
Answer by rfer(16322)

0.06x+0.08(10000-x)=760
0.06x+800-0.08x=760
-0.02x=-40
x=2000 at 6%
10000-x=8000 at 8%

Answer by mananth(16949)

Last year, Evelina invested $10,000.00, part at 6% annual interest and part at 8% annual interest. If she received $760.00 in interest at the end of the year, how much did she invest at each rate?
...
Total $10,000.00,
at 6% --------------$x
at 8% --------------$10000-x
annual interestreceived $760.00
...
6x+8(10,000-x)=760*100
..
6x+80,000-8x=76,000
-2x+80,000=76,000
-2x=-80,000+76,000
-2x=-4,000
/-2
x=$2000 @ 6%
$8,000 @ 8%
...
m.ananth@hotmail.ca